The young modulus of a material of total extension 1.2m if the sites produced on performing SHM is 20N/M^2 of 4m of diameter length the wire lenght of the sites is, 6.5m if a mass of50kg is suspended on that wire to causes it to osclillate what is the period of oscillation and the angular frequency

1
Expert's answer
2022-01-31T16:22:15-0500

The period of oscillation and the angular frequency:


"T=2\pi\sqrt{m/k}=9.9\text{ s},\\\space\\
\omega=\sqrt{k/m}=0.63\text{ rad/s}."